OpenStudy (anonymous):

Can someone answer this question for me please. i just need the answers... Solve. x^2 + 5x + 3 = 0

Parth (parthkohli):

Hi, welcome... but the site doesn't allow you to get answers. I will help you out if you need it. :-)

OpenStudy (hba):

You have to compare your equation with the standard form of quad equation and determine a,b and c. \[\huge\ ax^2+bc+c=0\] After you are done with that use the Quad formula, \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

Parth (parthkohli):

Heard about the quadratic formula? :-)

OpenStudy (hba):

There are two other methods you can use: 1)Breaking the middle term 2)Completing the square method.

OpenStudy (anonymous):

yes. but i don't understand how to solve this

OpenStudy (hba):

Do you see what all i have written up.

OpenStudy (anonymous):

i see it @ hba

OpenStudy (hba):

@keshab1234 So what is a,b and c ?

OpenStudy (hba):

1x^2 + 5x + 3 = 0 compare with ax^2+bx+c=0 a,b and c=

Parth (parthkohli):

I'd probably rephrase it, and ask: What are the coefficients?

OpenStudy (hba):

@ParthKohli Ye Americion ki dimaag mai kuch nahi jaata :/

OpenStudy (anonymous):

coefficients are substitutes for the actual number

OpenStudy (hba):

He asked what are the coefficients in your equation.

OpenStudy (amoodarya):